### v1

# BemaGANv2: A Tutorial and Comparative Survey of GAN-based Vocoders for Long-Term Audio Generation

Abstract
This paper presents a tutorial-style survey and implementation guide of BemaGANv2, an advanced GAN-based vocoder designed for high-fidelity and long-term audio generation. Built upon the original BemaGAN architecture, BemaGANv2 incorporates major architectural innovations by replacing traditional ResBlocks in the generator with the Anti-aliased Multi-Periodicity composition (AMP) module, which internally applies the Snake activation function to better model periodic structures. In the discriminator framework, we integrate the Multi-Envelope Discriminator (MED), a novel architecture we proposed, to extract rich temporal envelope features crucial for periodicity detection. Coupled with the Multi-Resolution Discriminator (MRD), this combination enables more accurate modeling of long-range dependencies in audio. We systematically evaluate various discriminator configurations, including MSD + MED, MSD + MRD, and MPD + MED + MRD, using objective metrics (FAD, SSIM, PLCC, MCD) and subjective evaluations (MOS, SMOS). This paper also provides a comprehensive tutorial on the model architecture, training methodology, and implementation to promote reproducibility. The code and pre-trained models are available at: https://github.com/dinhoitt/BemaGANv2.
keywords:
BemaGANv2 , Vocoder , Multi-Envelope Discriminator(MED) , Generative Adversarial Network (GAN) , Long-Term Audio Generation , High-Fidelity Audio , Periodicity Modeling
journal: ICT Express
1 Introduction
In recent years, deep learning-based models for Text-to-Audio (TTA) and Text-to-Music (TTM) Models have attracted considerable interest, in the context of AI-driven music Generation. One of the key components of diffusion-based generative models is the vocoder, which transforms intermediate acoustic representations such as Mel-spectrograms into raw waveform audio signals. The performance and characteristics of the vocoder significantly influence the realism and fidelity of the generated audio, particularly for long-duration outputs.
To address limitations in periodicity modeling and long term dependency, our research group previously introduced BemaGAN[1], a GAN-based vocoder inspired by the HiFi-GAN[2] framework. Notably, BemaGAN replaced the standard Multi-Scale Discriminator (MSD) with a custom-designed Multi-Envelope Discriminator (MED)[1, 3], enabling more sensitive detection of temporal envelope variations to enhance long-term audio quality.
Building upon this foundation, we developed BemaGANv2[3], an improved architecture that incorporates multiple architectural enhancements. These include the integration of the Anti-aliased Multi-Periodicity composition (AMP) module[4] in place of conventional ResBlocks within the generator, and the use of the Snake activation function[5] to strengthen periodic signal modeling. In addition, the frequency coverage was extended from 8kHz to 12kHz with a 24kHz sampling rate, and the discriminator was restructured by introducing the Multi-Resolution Discriminator (MRD) in place of the Multi-Period Discriminator (MPD).
This paper serves two primary purposes:
  • Provide a tutorial-style explanation of the architectural principles, training methodology, and implementation details of BemaGANv2.
  • Validate the effectiveness of the discriminator combination used in BemaGANv2 by conducting a comparative analysis of various discriminator configurations, including MSD+MED, MSD+MRD, and MPD+MED+MRD, and evaluate their performance using both objective and subjective metrics.
  • By consolidating the architectural insights and experimental outcomes, this paper aims to facilitate understanding and reproducibility of advanced vocoder systems, and to support further research on long-term, high-fidelity audio generation.
All experiments were conducted using the LJSpeech dataset for training and random audio samples from freesound.org for testing, on an NVIDIA A100 GPU environment via Google Colab.
This article is an extended version of our ICAIIC 2025 paper[3], enriched with additional experiments, theoretical and mathematical analysis, and presented as a survey and tutorial to enhance clarity and reproducibility.
2 Related Work
2.1 Overview of Vocoders
In TTA and TTM systems, the vocoder plays a critical role in converting intermediate acoustic representations—typically Mel-spectrograms—into time-domain audio waveforms. The quality and structure of the vocoder significantly affect the fidelity, naturalness, and temporal coherence of the generated audio. In fact, most diffusion-based TTA and TTM models, such as AudioLDM[6, 7],FLUX Music[8], commonly adopt HiFi-GAN as the vocoder due to its high-quality reconstruction and efficient inference. In this subsection, we review key developments in parallel GAN-based vocoders, focusing on MelGAN[9], HiFi-GAN, and BigVGAN, to illustrate the evolution of architecture design and its implications for periodicity modeling.
2.1.1 MelGAN
MelGAN was one of the earliest successful attempts to apply generative adversarial networks (GANs) to the task of non-autoregressive raw waveform generation. The model adopts a fully convolutional generator that transforms Mel-spectrograms into audio waveforms in a feed-forward and parallel manner, enabling real-time synthesis on both GPU and CPU without the need for autoregressive decoding or teacher-student distillation.
The generator in MelGAN employs transposed convolutional layers for upsampling, each followed by residual stacks with dilated convolutions to expand the receptive field and capture long-range dependencies. The design also avoids using global noise vectors, relying solely on the conditioning Mel-spectrogram as input. This was motivated by empirical findings that strong conditioning signals are sufficient to generate coherent audio without additional stochasticity.
For the discriminator, MelGAN introduces a multi-scale discriminator (MSD) architecture, where three sub-discriminators operate on different temporal resolutions of the input waveform. This design allows the discriminator to capture audio structures at varying scales, from local textures to global continuity[9].
Although MelGAN demonstrated impressive speed and efficiency—with more than 2500 kHz inference speed on a single GPU and a small parameter footprint ( 4.2M)—it showed perceptual quality limitations compared to autoregressive models like WaveNet or flow-based models like WaveGlow. Later studies, including HiFi-GAN, extended this line of work improved fidelity through enhanced discriminator designs and loss functions[3].
2.1.2 HiFi-GAN
HiFi-GAN is a high-fidelity and efficient GAN-based vocoder that significantly improved upon previous models like MelGAN by refining both the generator and the discriminator architectures. While it retains the Multi-Scale Discriminator (MSD) introduced in MelGAN, HiFi-GAN adds a novel Multi-Period Discriminator (MPD) to more effectively capture the periodic structures of waveform signals[3].
The generator in HiFi-GAN adopts a stack of upsampling layers followed by multi-receptive field fusion (MRF) modules, which consist of multiple residual blocks with varying kernel sizes and dilation rates. This design enables the generator to capture both local and global patterns in the signal, contributing to better harmonic reconstruction[9].
The Multi-Period Discriminator (MPD) is the key innovation in HiFi-GAN’s discriminator framework. Each sub-discriminator in the Multi-Period Discriminator (MPD) reshapes the 1-D waveform into 2-D matrices using different periodic lengths and applies 2-D convolutions to capture periodic patterns over varying cycles. This architecture complements the Multi-Scale Discriminator (MSD), which focuses on multi-scale representations in the time domain via downsampling[9].
HiFi-GAN achieves high fidelity in waveform generation while maintaining extremely fast inference speed (e.g., 93.75× faster than real-time on GPU), making it suitable for real-time TTS systems. However, its reliance on Leaky ReLU activation functions and conventional residual blocks limits its capacity to learn fine-grained periodicity and harmonic richness in more complex or out-of-distribution (OOD) data, as pointed out in subsequent works like BigVGAN[5].
Table 1: Comparison of GAN-based vocoder architectures.
Model Generator Discriminator(s) Activation Key Features
MelGAN Transposed Conv + ResBlock MSD Leaky ReLU Fast inference, low fidelity
HiFi-GAN MRF-based ResBlocks MPD + MSD Leaky ReLU Harmonic modeling, MRF design
BigVGAN AMP block MPD + MRD Snake Anti-aliasing, large-scale training
BemaGAN HiFi-GAN-style ResBlock MED + MPD Leaky ReLU Envelope-aware discrimination
BemaGANv2 AMP block MED + MRD Snake Balanced fidelity and periodicity
As a result, HiFi-GAN has served as a strong baseline for later vocoders aiming to improve both periodicity modeling and generalization, including BigVGAN and the BemaGAN series.
2.1.3 BigVGAN
BigVGAN is a universal GAN-based vocoder designed to synthesize high-fidelity audio across a wide range of conditions, including unseen speakers, languages, and noisy recording environments. It builds upon the generator architecture of HiFi-GAN but introduces several architectural and training innovations to address the limitations of prior models in capturing periodicity, extrapolation, and robustness to out-of-distribution (OOD) inputs[5].
The core architectural contribution of BigVGAN is the AMP (Anti-aliased Multi-Periodicity) block, which replaces the traditional residual blocks in the generator. Each AMP block consists of a series of dilated convolutions embedded with channel-wise Snake activation functions, thereby enforcing an explicit periodic inductive bias in the transformation process. Each AMP block contains a series of dilated convolutions with channel-wise Snake activation functions, which provide an explicit periodic inductive bias. The Snake function, defined as:
$f_{\alpha}(x)=x+\frac{1}{\alpha}\sin^{2}(\alpha x)$ (1)
injects learnable periodicity into the activation, enabling the generator to better model complex harmonic structures in audio. This significantly improves the model’s ability to generalize and extrapolate periodic content, which is particularly beneficial in musical or expressive vocal signals[4].
To prevent aliasing artifacts caused by high-frequency nonlinearities, BigVGAN applies a low-pass filter to the output of the Snake activation via a process of anti-aliased upsampling and downsampling, inspired by alias-free design principles from image generation models like StyleGAN3[4, 10].
In the discriminator, BigVGAN replaces HiFi-GAN’s Multi-Scale Discriminator (MSD) with a Multi-Resolution Discriminator (MRD), which operates on spectrograms of various STFT configurations. This Multi-Resolution Discriminator (MRD), originally proposed in UnivNet, enables enhanced spectral sharpness and pitch accuracy by leveraging multiple time-frequency resolutions[11].
BigVGAN is also notable for its large-scale training: the full version of the model contains over 112 million parameters and is trained on the entire LibriTTS corpus, including diverse environments. Despite the increased capacity, it achieves faster-than-real-time inference and state-of-the-art results on both in-distribution and out-of-distribution tasks, outperforming HiFi-GAN and other vocoders in both MOS and SMOS metrics.
In summary, BigVGAN demonstrates that enhancing the generator’s inductive bias through periodic activations and anti-aliasing, along with robust discriminator design and scaling strategies, leads to significant improvements in universal vocoding performance. However, the increased architectural complexity and sensitivity during training highlight the need for more efficient and stable vocoder designs in future work.
2.2 Snake Function and Periodictiy in Audio
This limitation is closely tied to the inductive biases of common activation functions. Functions like ReLU and its variants extrapolate linearly, while tanh and sigmoid tend to saturate to constants. As a result, these activations are inherently ill-suited for learning or generating periodic signals, which require oscillatory behavior that persists outside the training interval.
To address this, recent research has proposed the Snake activation function, which introduces a trainable periodic component while maintaining monotonicity and optimization stability. Unlike earlier periodic activations such as sin or cos—which are difficult to optimize due to non-monotonicity and numerous local minima—Snake balances periodicity and trainability[5].
Theoretically, the Snake activation satisfies a universal extrapolation property: neural networks using Snake can approximate any piecewise-smooth periodic function over the real line, not just within the training interval. Empirically, it has been shown to outperform ReLU, tanh, and sin in periodic regression tasks and time-series prediction.
In the context of neural vocoders, Snake provides a strong inductive bias for harmonic structure, enabling better modeling of voiced speech and musical tones. When embedded within advanced generator architectures, it enhances both periodicity modeling and signal continuity, making it a powerful component for high-fidelity audio synthesis.
2.3 Architectural Comparisons and Motivations for BemaGANv2
While previous sections have detailed individual vocoder architectures, we here provide a comparative analysis focusing on their strengths, limitations, and how they motivated the design of BemaGANv2.
HiFi-GAN combines Multi-Scale Discriminator (MSD) and Multi-Period Discriminator (MPD) with a residual-based generator, enabling high-fidelity audio synthesis[2]. However, its generator lacks a strong inductive bias for periodicity, and the fixed-period hop sizes used in Multi-Period Discriminator (MPD) can be suboptimal for diverse or expressive signals.
BigVGAN improves upon this by introducing the AMP block with periodic activations (Snake) and replacing Multi-Scale Discriminator (MSD) with the Multi-Resolution Discriminator (MRD) for enhanced spectral sharpness[4]. However, this comes with a significant increase in parameter count and training complexity.
BemaGAN seeks to improve temporal dynamics modeling through the Multi-Envelope Discriminator (MED), but retains generator of HiFi-GAN, which limits the benefit of the improved discriminator[1].
BemaGANv2 is thus designed to combine the periodicity modeling capability of the AMP module (used in BigVGAN) with the temporal envelope sensitivity of Multi-Envelope Discriminator (MED) and the spectral resolution of Multi-Resolution Discriminator (MRD). This MED+MRD combination enables better coverage of both time-domain and frequency-domain perceptual cues while keeping the model more lightweight and stable than BigVGAN.
This integration addresses the limitations of each predecessor and is empirically validated in Section 4.
3 Model
In this section, we provide a detailed tutorial-style description of the architectural components of BemaGANv2. The model adopts a GAN-based vocoder framework, integrating periodicity-aware generator and perceptually motivated discriminator. We begin by describing the generator based on AMP blocks, followed by the Multi-Envelope Discriminator (MED) and Multi-Resolution Discriminator (MRD), and conclude with the training objectives.
3.1 Generator with AMP Block
Figure 1: The architecture of the AMP-based generator used in BemaGANv2. The AMP block, originally introduced in BigVGAN, integrates upsampling and downsampling operations with low-pass filtering (LPF), Snake activation for periodic inductive bias, and dilated convolutions.
The generator in BemaGANv2 is structured similarly to HiFi-GAN, using transposed convolution layers for upsampling and residual blocks for waveform synthesis. However, instead of conventional residual blocks, it employs the AMP block, originally introduced in BigVGAN[4].
Each AMP block contains dilated convolutions paired with the Snake activation function, which provides a learnable periodic bias. This allows the generator to better model harmonic structures and rhythmic patterns in long-term audio generation. The overall design enhances both local detail and global periodic consistency[4][5].
3.2 Multi-Envelope Discriminator (MED)
Figure 2: The structure of the Multi-Envelope Discriminator (MED). Time-domain envelopes, including both upper and lower envelopes, are extracted from the input audio using low-pass filters with different cutoff frequencies. These envelope signals are then processed by 1D convolutional layers. This design enables the discriminator to detect temporal energy patterns that are crucial for perceptual quality, such as prosodic variation.
A central contribution of BemaGANv2 is the introduction of the Multi-Envelope Discriminator (MED)[1][3], a novel discriminator architecture proposed by our research group. Rather than operating directly on raw waveforms or spectrograms, Multi-Envelope Discriminator (MED) analyzes time-domain envelopes derived from the audio signal. These envelopes are extracted using multiple low-pass filters with varying cutoff frequencies, capturing energy dynamics at different temporal scales.
Each envelope is processed by a dedicated 1D convolutional stack, and the outputs are aggregated to form the discriminator output. This design enables the model to detect patterns related to phrasing, prosody, and amplitude modulation properties closely linked to human perception of naturalness. Multi-Envelope Discriminator (MED) thus plays a crucial role in enhancing long-term temporal coherence in the generated audio.
3.3 Multi-Resolution Discriminator (MRD)
To complement Multi-Envelope Discriminator (MED), BemaGANv2 also incorporates a Multi-Resolution Discriminator (MRD). This module operates in the time-frequency domain using STFT-based log-magnitude spectrograms. By processing multiple spectrograms computed with different FFT sizes and hop lengths, Multi-Resolution Discriminator (MRD) enforces spectral consistency across various resolutions.
The Multi-Resolution Discriminator (MRD) primarily helps the model maintain harmonic sharpness and timbral accuracy, which are important for high-fidelity generation. Together with Multi-Envelope Discriminator (MED), Multi-Resolution Discriminator (MRD) provides complementary perspectives temporal envelope and spectral structure guiding the generator to produce perceptually convincing waveforms.
3.4 Overall Architecture
Figure 3 presents the complete BemaGANv2 architecture. The generator transforms the input Mel-spectrogram into a waveform using AMP blocks. The output waveform is then assessed by both Multi-Envelope Discriminator (MED) and Multi-Resolution Discriminator (MRD), which provide feedback based on temporal energy flow and spectral consistency, respectively. This dual-discriminator setup enables BemaGANv2 to generate audio that is both rhythmically natural and spectrally accurate.
Figure 3: Overview of the BemaGANv2 architecture. The generator converts Mel-spectrograms into raw waveforms, which are evaluated by two discriminators: the Multi-Envelope Discriminator (MED) and the Multi-Resolution Discriminator (MRD). Adversarial and auxiliary losses are backpropagated from both discriminators to train the generator.
3.5 Training Loss
The training loss in BemaGANv2 follows the strategy proposed in HiFi-GAN and is based on the LSGAN formulation. Unlike the original GAN objective, which uses binary cross-entropy, LSGAN replaces it with a least squares loss to improve training stability and maintain non-vanishing gradients[12].
Adversarial Loss: Let $D$ denote the discriminator and $G$ the generator. The discriminator is trained to assign real samples a value close to 1 and generated samples a value close to 0. The generator is optimized so that the generated samples fool the discriminator, i.e., are classified close to 1.
$\mathcal{L}_{Adv}(D;G)=\mathbb{E}_{x,s}[(D(x)-1)^{2}+(D(G(s)))^{2}]$ (2)
$\mathcal{L}_{Adv}(G;D)=\mathbb{E}_{s}[(D(G(s))-1)^{2}]$ (3)
Mel-Spectrogram Loss: We adopt an $L_{1}$ loss between the Mel-spectrograms of the real and generated waveforms:
$\mathcal{L}_{Mel}(G)=\mathbb{E}_{x,s}[\|\phi(x)-\phi(G(s))\|_{1}]$ (4)
where $\phi(\cdot)$ is the Mel-spectrogram transformation.
Feature Matching Loss: This loss minimizes the $L_{1}$ distance between intermediate features of the discriminator for real and generated samples:
$\mathcal{L}_{FM}(G;D)=\sum_{l=1}^{L}\mathbb{E}[\|D^{(l)}(x)-D^{(l)}(G(s))\|_{1}]$ (5)
Total Losses: The final loss functions for the generator and discriminator are given as:
$\mathcal{L}_{G}=\sum_{k=1}^{K}[\mathcal{L}_{Adv}(G;D_{k})+\lambda_{fm}\mathcal{L}_{FM}(G;D_{k})]+\lambda_{mel}\mathcal{L}_{Mel}(G)$ (6)
$\mathcal{L}_{D}=\sum_{k=1}^{K}\mathcal{L}_{Adv}(D_{k};G)$ (7)
where $K$ is the number of sub-discriminators, and $\lambda_{fm},\lambda_{mel}$ are hyperparameters for feature matching and mel-spectrogram losses, respectively.
4 Experiments and Results
In this section, we conduct a comprehensive evaluation of various discriminator configurations and their impact on audio reconstruction quality. We place particular emphasis on the architectural combinations used in BemaGANv2, while also referencing models such as HiFi-GAN, BigVGAN, and our previous BemaGAN. The goal is not only to compare model performance, but also to offer practical insights for vocoder design choices in short-term and long-term audio synthesis tasks. For short-term evaluation, we used audio clips of approximately 20 seconds, including a diverse range of content such as human speech, drum and percussion sounds, short musical phrases, and car engine noise. For long-term evaluation, we used 90-second segments of full music tracks spanning various genres. To ensure a fair comparison of discriminator configurations, all models except HiFi-GAN employed a generator architecture based on AMP blocks with embedded Snake activation functions.
4.1 Evaluation Metrics
To ensure both analytical rigor and perceptual validity, we adopt four objective and two subjective metrics:
Objective Metrics
  • Fréchet Audio Distance (FAD): Quantifies the distributional divergence between the generated and the ground-truth audio embeddings[13].
  • Structural Similarity Index (SSIM): Evaluates the perceptual closeness between Mel-spectrogram images across RGB channels[14].
  • Pearson Correlation Coefficient (PLCC): Measures the degree of linear correlation in frequency domain characteristics[15].
  • Mel-Cepstral Distortion (MCD): Captures the distortion in MFCC sequences using Dynamic Time Warping[16].
In our implementation, we adopt the standard SPTK-based configuration and compute MCD using fastDTW after extracting the mel-cepstrum from each frame. While the reported MCD values in our table may exceed 1 due to the evaluation on relatively long-term (30s, 90s) audio and due to practical constraints, such as limited computing resources, short training epochs, and the need to evaluate over six different model configurations. However, when we test under typical conditions used in prior work—namely short-form utterances of around 10 seconds and sufficiently trained models—we are able to achieve MCD values well below 1.0, which we have confirmed empirically. This validates the effectiveness and correctness of our MCD implementation.
Subjective Metrics
  • MOS (Mean Opinion Score): Participants rate the overall perceptual quality of the audio samples without seeing the ground truth.
  • SMOS (Similarity MOS): Participants compare the similarity between the generated samples and the reference ground-truth audio.
All subjective evaluations are conducted on a 5-point scale with 95% confidence intervals, using samples from both short- and long-term audio.
4.2 Objective Evaluation Results
Table 2: Short-Term Audio Objective Metrics
Models FAD $\downarrow$ SSIM $\uparrow$ PLCC $\sim 1$ MCD $\downarrow$
BemaGANv2 0.911 0.78 0.949 1.652
BemaGAN 3.194 0.70 0.861 3.18
BigVGAN 1.906 0.71 0.906 2.56
HiFi-GAN 6.06 0.73 0.792 3.22
MSD + MED 2.337 0.68 0.879 2.964
MSD + MRD 3.327 0.66 0.841 3.354
MED + MPD + MRD 2.146 0.71 0.851 3.136
Table 3: Long-Term Audio Objective Metrics
Models FAD $\downarrow$ SSIM $\uparrow$ PLCC $\sim 1$ MCD $\downarrow$
BemaGANv2 2.685 0.39 0.945 1.8
BemaGAN 6.612 0.43 0.866 2.841
BigVGAN 3.58 0.35 0.908 2.28
HiFi-GAN 30.883 0.53 0.702 7.75
MSD + MED 6.66 0.35 0.810 6.92
MSD + MRD 18.27 0.39 0.776 7.52
MED + MPD + MRD 5.969 0.40 0.878 2.7557
While our earlier experiments indicated that HiFi-GAN performed better in short-duration tasks, additional training with extended epochs has revealed a different trend[3]. BemaGANv2, when trained for a longer duration, not only maintains its superior performance in long-term audio generation but also clearly outperforms other models in short-term tasks.
We hypothesize that this improvement is linked to the gradient behavior of the Snake activation function. Unlike Leaky ReLU, which maintains a non-zero gradient across its domain, Snake contains regions where the derivative approaches zero, potentially slowing convergence in early training stages. However, with sufficient training, the model leverages the periodic inductive bias of Snake to capture harmonic and transient features more effectively.
A detailed mathematical explanation of this behavior, including derivative analysis, is provided in  A.
4.3 Subjective Evaluation Results
Table 4 and Table 5 report the results of human evaluations, which are broadly consistent with the objective measurements. These results provide further evidence that architectural decisions—such as the choice of activation functions and discriminator types—impact the perceptual quality across different audio durations.
Table 4: Mean Opinion Score (MOS)
Models Short Audio MOS $\uparrow$ Long Audio MOS $\uparrow$
Ground Truth 3.67(±0.09) 4.73(±0.12)
BemaGANv2 3.08(±0.10) 3.46(±0.12)
BemaGAN 1.74(±0.11) 2.67(±0.13)
BigVGAN 2.52(±0.14) 3.07(±0.13)
HiFi-GAN 2.91(±0.12) 1.14(±0.09)
MSD + MED 2.16(±0.09) 2.313(±0.14)
MSD + MRD 2.14(±0.19) 2.37(±0.15)
MED + MPD + MRD 2.17(±0.09) 2.31(±0.12)
Table 5: Similarity Mean Opinion Score (SMOS)
Models Short Audio SMOS $\uparrow$ Long Audio SMOS $\uparrow$
BemaGANv2 3.07(±0.09 3.53(±0.09)
BigVGAN 2.61(±0.12) 3.09(±0.09)
HiFi-GAN 2.66(±0.12) 1.03(±0.03)
Consistent with the objective evaluation results, BemaGANv2 received the highest subjective scores in both perceptual quality (MOS) and similarity (SMOS), further validating its effectiveness across different audio durations.
4.4 Comparing Mel-spectrogram Structure
The visualization of the Mel spectrogram in Figure 4 reinforces these findings. BemaGANv2 shows better structure preservation and noise suppression in long-term audio compared to BigVGAN.
Figure 4: Mel-Spectrogram visualization of samples from Ground Truth, BigVGAN, and BemaGANv2 trained on LJSpeech.
4.5 Result
The results clearly demonstrate that BemaGANv2 achieves superior performance in both short- and long-term audio tasks, outperforming all other models in objective and subjective evaluations. This highlights the effectiveness of its generator design and discriminator configuration, particularly in preserving both high-frequency details and long-term periodic structures.
Although BemaGANv2 shows relatively low SSIM scores, this can be attributed to its behavior in reconstructing amplitude levels. SSIM is computed by comparing Mel-spectrogram images across three RGB channels, where the color intensity primarily reflects the sound loudness. The lower SSIM indicates that BemaGANv2 tends to deviate from the exact decibel scale of the original audio. However, subjective metrics such as MOS and SMOS consistently rank BemaGANv2 as the top performer. This suggests that human listeners are more sensitive to the perceptual fidelity of the audio, such as timbral and structural quality, than to precise restoration of loudness. In support of this, BemaGANv2 also achieves the best scores in FAD and MCD, which are directly related to perceptual realism and spectral coherence.
Interestingly, although the MED+MPD+MRD combination yielded decent objective metrics — especially in long-term evaluation — it showed significantly lower subjective scores. Through an analysis of training and validation loss curves, we confirmed that the issue is not due to overfitting. Instead, we attribute this degradation in perceptual quality to mode collapse[17], likely caused by adversarial instability when employing an excessive number of discriminators. Following this observation, we did not pursue further experiments involving combinations of three or more discriminators, as the risk of mode collapse outweighed potential performance gains.
These findings emphasize the importance of balancing feedback diversity and adversarial stability in discriminator design. For further experimental details, including loss curves and configuration analysis, please refer to  B.
5 Discussion
During the evaluation of vocoder outputs, we observed a peculiar issue in the HiFi-GAN model where long-term audio samples resulted in outputs nearly double the expected length. This phenomenon was not observed in other models (e.g., BigVGAN, BemaGANv2), despite identical training data, architectures, and preprocessing pipelines.
To identify the source of this anomaly, we systematically investigated several components of the inference pipeline:
  • Configuration consistency between training and inference (e.g., hop_size, n_fft) — confirmed to be consistent.
  • STFT and Mel-spectrogram center alignment — identical across models and not responsible for the anomaly.
  • Dataset bias — all models were trained on LJSpeech, which consists of short, speech-based audio samples.
  • Model architecture — no differences except the activation function.
The only significant difference between HiFi-GAN and the other models lies in the activation function: HiFi-GAN utilizes LeakyReLU[2], while others incorporate the Snake function. Given that Snake is known to better extrapolate periodic structures, especially in long audio sequences[5], we conclude that the root cause of the abnormal doubling in waveform duration stems from HiFi-GAN’s inability to generalize beyond the training distribution of short-length signals.
This result underscores the critical role of periodic inductive bias in vocoders tasked with long-term synthesis. Snake-based models consistently exhibit better stability and fidelity.
For reproducibility and further inspection, debugging logs and comparative duration graphs are provided in the  C
6 Conclusion
In this study, we presented BemaGANv2, a vocoder designed to enhance both the fidelity and temporal coherence of waveform reconstruction. Through theoretical motivations and extensive empirical evaluations, we demonstrated that BemaGANv2 outperforms existing GAN-based vocoders such as HiFi-GAN and BigVGAN, particularly in long-duration audio tasks. The model’s generator architecture incorporating AMP blocks and the proposed Multi-Envelope Discriminator (MED) proved effective in capturing both local details and long-term periodicity.
Our results showed that BemaGANv2 achieves superior performance in both objective and subjective evaluation metrics in short- and long-term audio. Notably, it exhibited perceptual advantages even in scenarios where traditional metrics such as SSIM were lower, highlighting its ability to preserve the auditory quality that human listeners prioritize.
These findings collectively suggest that BemaGANv2 is a highly suitable vocoder for Text-to-Music (TTM) and Text-to-Audio (TTA) models, especially diffusion based generation pipelines. As a next step, we plan to integrate BemaGANv2 into multimodal generative AI systems, enabling more expressive and temporally coherent audio generation conditioned on various modalities. We believe this direction will further bridge the gap between machine-generated and human-perceived audio realism.
Acknowledgments
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00352526)
Appendix A Analysis of the Activation Function
Our earlier experiments suggested that HiFi-GAN performs best on short-duration audio[3], especially under limited training epochs. However, in the extended study presented in this paper, we provided sufficient training epochs, and the results indicate that BemaGANv2 outperforms all baselines in both short- and long-term waveform reconstruction tasks. In this section, we provide a mathematical interpretation to explain this behavior, particularly in relation to the activation functions used.
A.1 Periodic Bias of the Snake Function
Snake function is designed to introduce periodic inductive bias, which makes it highly effective at capturing repeating structures in long-term audio signals. However, in short audio segments—especially those dominated by non-periodic transient features or high-frequency details—this periodic bias may cause the model to underfit rapid variations, leading to degraded performance if not sufficiently trained.
A.2 Gradient Vanishing Issue
The derivative of the Snake function is:
$f_{\alpha}^{\prime}(x)=1+\sin(2\alpha x)$ (A.1)
When $\sin(2\alpha x)=-1$, this derivative periodically approaches zero. In such regions, the gradient flow is suppressed, making it difficult for the model to update its weights during training. This is especially problematic in learning small amplitude variations common in short-term signals, where precise gradient flow is crucial.
In contrast, Leaky ReLU[19] is defined as:
$\mathrm{LeakyReLU}(x)=\begin{cases}x,&\text{if }x\geq 0\\ \beta x,&\text{otherwise}\end{cases}$ (A.2)
and its derivative is constant:
$\mathrm{LeakyReLU}^{\prime}(x)=\begin{cases}1,&\text{if }x\geq 0\\ \beta,&\text{otherwise}\end{cases}$ (A.3)
This consistent gradient ensures stable learning even in shallow networks or early training stages. Therefore, without enough training iterations, Snake-based generators may struggle to achieve the same level of short-term fidelity as those using Leaky ReLU. Our results suggest that when given sufficient epochs, the advantages of Snake’s periodic bias are fully realized, enabling BemaGANv2 to excel in both short- and long-term tasks.
Appendix B Loss Curve
Figure B.1: Training loss curves of BemaGANv2, HiFi-GAN, and BigVGAN.
Figure B.2: Training and validation loss curves of the MED+MPD+MRD combination.
Figure B.1 presents the training loss curves for BemaGANv2, HiFi-GAN, and BigVGAN. The results indicate that BemaGANv2 demonstrates superior training stability compared to the other models. This suggests that the combination of Multi-Envelope Discriminator (MED) and Multi-Resolution Discriminator (MRD) discriminators in BemaGANv2 provides more consistent adversarial feedback and contributes to robust convergence.
Figure B.2 illustrates the training and validation loss curves for the MED+MPD+MRD configuration. Although previous studies raised concerns about potential overfitting in multi-discriminator setups, the curve clearly shows that the observed performance drop in subjective metrics is not due to overfitting. Rather, the consistency between training and validation curves supports our conclusion that the degradation is caused by mode collapse resulting from adversarial instability.
Appendix C Debugging Logs and Analysis
To diagnose the anomalous behavior observed during inference with HiFi-GAN—where long-term audio samples were generated at nearly double the intended length—we conducted a series of logging and validation procedures. This appendix provides sample logs and graphical summaries of the debugging process.
C.1 Waveform Length Validation Log
The following logs show the comparison between original and generated waveform lengths for multiple test samples. All lengths are computed in samples and seconds.
===== Inference: example_1.wav =====
Original sampling rate: 44100
Original waveform length: 4,279,739 samples
Original duration: 97.05 sec
Mel shape: torch.Size([1, 80, 16717])
Expected waveform length: 4,279,552 samples
Generated waveform length: 4,279,552 samples
Length difference: 0 samples (0.00 sec)
===== Inference: example_2.wav =====
Original waveform length: 3,873,082 samples
Generated waveform length: 3,873,024 samples
Difference: 0 samples (0.00 sec)
C.2 Length Consistency Visualization
Figure C.1: Comparison between original and generated waveform lengths for short and long audio inputs.
Figure C.1 confirms that waveform length discrepancies are not consistently present across all durations. Short clips (e.g., 20 seconds) are reconstructed with accurate length, while anomalies are observed predominantly in long-form audio in specific configurations.
C.3 Mel-to-Audio Consistency Verification
To verify the alignment between mel-spectrogram frame count and audio frame count, we tracked the following:
[…]
  • All inference samples conformed to this relationship.
C.4 Conclusion
After validating inference logs, comparing mel frame counts, and matching waveform durations, we ruled out issues in pre-processing, STFT padding, or inference time truncation. These observations suggest that the previously reported duration anomaly is not caused by implementation bugs, but likely arises from training data mismatch, activational bias (e.g., LeakyReLU vs. Snake), or temporal extrapolation failure specific to HiFi-GAN under long-term generation settings.
Conflict of interest
The authors declare that there is no conflict of interest in this paper.
Declaration of Generative AI and AI-assisted Technologies in the Writing Process
During the preparation of this work, the author(s) used ChatGPT in order to improve English translation and language clarity. After using this tool, the author(s) reviewed and edited the content and take full responsibility for it.